OpenStudy (anonymous):

Which is the equation to the line passing through points (2, –7) and (5, –1)

OpenStudy (raden):

use the formula : (y-y1)/(y2-y1) = (x-x1)/(x2-x1)

OpenStudy (anonymous):

use the formula @RadEn gave you to find the slope of your line first.

OpenStudy (anonymous):

y = 2x – 8 y = – 2x + 9 y = 2x – 11 y = – 2x + 6

OpenStudy (anonymous):

@Brittni0605 those are the answers i think its A

OpenStudy (anonymous):

No, its not A.

OpenStudy (anonymous):

then at @Brittni0605 what is it?

OpenStudy (anonymous):

Show me work for finding the slope of your line given the two points in your problem

OpenStudy (anonymous):

dont know how to

OpenStudy (anonymous):



OpenStudy (anonymous):

Understand?

OpenStudy (anonymous):

somewhat so does that mean that the answer would be c ?

OpenStudy (anonymous):

It means the answer could be A or C.

OpenStudy (anonymous):

But since I already told you its not A, you know its C.

OpenStudy (anonymous):

i asked u last time if it was A & you told me no its not

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

I said its not A. But based on just slope, your choices were narrowed down to A or C. To determine if it was A or C, you needed to find the b value in the equation y=mx+b